Solve (x+y)(p+q) =z-1
1
Expert's answer
2020-07-01T18:21:38-0400

Given "(x+y)(p+q) =z-1"

"\implies (x+y)p + (x+y)q = z-1"

So, Lagrange's equation is

"\frac{dx}{x+y} = \frac{dy}{x+y} = \frac{dz}{z-1}" .

By First two factors of Lagrange's equation

"\frac{dx}{x+y} = \frac{dy}{x+y} 
\\
\implies dx = dy \\
\implies x - y =c_1" ______________(1)

By last two factors, we get

"\frac{dy}{x+y} = \frac{dz}{z-1}"

Now, by using solution (1), we get

"\frac{dy}{c_1+2y} = \frac{dz}{z-1}"

By integrating, we get

"\frac{1}{2}\log(c_1+2y) = \log(z-1)+c_2"

"\implies \frac{1}{2}\log(x+y) = \log(z-1)+c_2"

"\implies 0 = \log(\frac{z-1}{\sqrt{x+y}})+c_2"

"\implies \frac{z-1}{\sqrt{x+y}} = e^{-c_2} = c_3" ____________(2)

Hence, Solution is "c_3 = f(c_1)"

"\implies z-1 = (\sqrt{x+y}) \ f(x-y)"